THE IONIZATION ENERGY OF CYANO PHOSPHINOGEN (HPCN)

\underline{PATRICK E. FLEMING}, Keck Facility for Chemical Research, Department of Chemistry, San Jose State University, San Jose, CA 95192.

Optimized structures and harmonic vibrational frequencies have been determined for cyano phosphinogen (HPCN) and the corresponding singly charged cation at the B3LYP/6-311++g(2df,p) level of theory. Total energies have also been determined at the G2 level of theory in order to predict the ionization energy of HPCN. In the ground state of HPCN^{+}, the molecule is bent (C_{s} symmetry) and the electronic state is of ^{1}A′ symmetry. The lowest-lying triplet state of the cation (^{3}A′′) lies approximately 1 eV above the singlet state and the molecule is also bent in this state (C_{s} symmetry). The energy ordering of states in the cation is unexpected by analogy to HNCN^{+} whose ground state is of triplet spin degeneracy. Also, the triplet state of HNCN^{+} is linear, unlike HPCN^{+}. Structures, harmonic frequencies and ionization energies will be presented for the cyano phosphinogen system and discussed in comparison the cyano amidogen system.